# Beyond Data Privacy: New Privacy Risks for Large Language Models

### v1

@@ 1 Introduction
Beyond the risks of training data leakage, privacy threats also emerge from the integration of LLMs into larger, more complex systems, which we refer to as LLM-powered systems. These systems, especially those that use LLMs as decision-making engines in agent-based applications [36, 169], introduce new vulnerabilities and expand the potential attack surface for privacy violations. For instance, a user may share personal information with an LLM-based chatbot in order to receive personalized responses or suggestions. However, this information could be exfiltrated through side channels [23] or unintentionally disclosed by the model itself [158]. Such risks are not inherent to the LLM alone but emerge from the architecture of interactions between users, models, and other system components. As LLM-powered applications become increasingly widespread in both daily life and professional domains, these privacy risks become more prominent and urgent.
A third category of privacy threats arises from the advanced reasoning and autonomous decision-making capabilities of LLMs, which create new opportunities for malicious exploitation. These capabilities enable adversaries to automate sophisticated attacks at unprecedented scale and speed, substantially lowering the barrier to entry for cyberattacks. For instance, an attacker could instruct an LLM to infer sensitive attributes, such as a user’s demographics, from their public online posts, leading to de-anonymization and other severe cybercrimes [49, 101]. Similarly, LLMs can be leveraged to launch large-scale, highly personalized social engineering campaigns, resulting in significant financial and societal consequences [106].
@@ 3.1 Membership Inference Attack
Closely connected with Differential Privacy (DP) [54, 95], Membership Inference Attack (MIA) [152] has become a widely adopted approach for privacy auditing of machine learning (ML) models [119, 157]. MIA assesses how much a trained ML model reveals about its training data by determining whether specific query instances (or documents) were included in the dataset. In the context of LLMs, membership inference attacks have been applied at different stages to explore the privacy risk.
@@
Potential Mitigation Strategies. Machine learning with differential privacy (e.g., DP-SGD [8] and PATE [131]) is an effective defense mechanism against privacy attacks, including MIAs. Studies [98, 135, 191] have applied DP-SGD to fine-tune LLMs on sensitive domains, which may degrade model utility under reasonable privacy budgets. For in-context learning scenarios, various DP-based techniques have been proposed, such as differential private prompt tuning [50, 70] and differential private synthetic text generation [179, 161]. In addition to the theoretical privacy guarantees that DP offers, many empirical privacy methods [199, 71] have shown effectiveness against MIAs. For example, LoRA [71], a widely used efficient fine-tuning method for LLMs, demonstrates better privacy preservation than full fine-tuning. However, these empirical defenses may be compromised when facing stronger MIAs [65], and there remains a notable trade-off between privacy and utility, especially for pre-trained LLMs.
@@ 4 Privacy Risks in LLM-Powered Systems
The integration of LLMs as core components in larger, complex systems introduces new vectors for privacy risks. For example, LLM-based chatbots, such as ChatGPT [127] and Claude [14], have become the primary interface through which users interact with LLMs. In these interactions, users often share personal narratives and sensitive details to seek advice or obtain personalized responses. This turns their conversation histories into a rich repository of private information, including personal preferences, habits, and even users’ secrets [114, 18, 85]. Like other computer systems, these applications are vulnerable to side-channel attacks, where adversaries exploit indirect information leaks from the system to steal data. Furthermore, the unique features of LLM-based applications, such as reasoning and memory mechanisms, provide additional attack surfaces. Below, we detail two prominent threats associated with LLM-powered systems: side channel attacks and information exfiltration.
@@ 4.2 Information Exfiltration
  • Unintended Disclosure. LLMs often lack awareness of privacy norms and the contextual boundaries of information flows [124]. As a result, they may inadvertently disclose sensitive information to inappropriate recipients. An LLM agent involved in a multi-round conversation may unintentionally repeat or expose previously shared user information, even when it is contextually irrelevant [115, 38, 148, 204]. For instance, it is undesirable for an LLM assistant to reveal that “John is talking to a few companies about switching jobs” while drafting an email to John’s current manager, particularly without his consent. This risk increases when LLMs are tasked with complex operations that involve integrating multiple sources of user data, such as combining financial, location, and preference information for personalized recommendations [148]. LLMs struggle to track which information is appropriate to share, which makes these disclosures particularly insidious.
@@
  • Leakge via Share Link. Users may share their conversations with a chatbot (e.g., ChatGPT) through a built-in “Share” button on commercial chatbot platforms, allowing only those with the link to access the conversation. However, it has been shown [6] that these links can be discovered through search engines (e.g., Google Search). This can unintentionally expose users’ conversation histories to the public. Furthermore, deleting a conversation from your ChatGPT history does not remove the public share link or prevent it from appearing in search engine results. Recognizing the potential for privacy breaches, OpenAI has addressed this issue by providing users with the option to control whether their chats are visible in search engine results, offering more control over users’ privacy.
@@
Vulnerability Detection and Mitigation Strategies. To counter the threat of information exfiltration, research efforts are focused on both detecting vulnerabilities and developing active defenses. On the detection front, some works measure an LLM’s capacity for privacy reasoning in ambiguous contexts to identify risks of unintended information disclosure [190, 149]. Another approach [16, 206] examines whether agents interacting with web interfaces adhere to the principle of data minimization, introducing benchmarks to systematically evaluate their compliance. In parallel, other efforts aim to build direct defenses against malicious attacks. This includes designing robust countermeasures for prompt injection attacks, which are a primary vector for information exfiltration [41, 168]. Despite these advancements, a comprehensive mitigation strategy capable of defending against the full spectrum of emerging exfiltration threats remains an open challenge [67, 200].
@@ 5.1 Automated Profile Inference
Individuals constantly generate digital footprints through their online activities, encompassing activities from social media comments and posts to shared photos and videos. While some of this data is inherently private (e.g., browse history), a vast amount of these activities (e.g., posts and comments) is publicly accessible. However, the public availability of this information does not eliminate privacy risks. An adversary can aggregate these seemingly innocuous public activities to construct a detailed personal profile, a process known as profiling [20, 47]. For instance, analyzing a Reddit user’s most frequented subreddits could reveal their hobbies, while geotags in posted images could disclose their travel patterns or home location. Profiling is widely recognized as a privacy violation by legitimate privacy frameworks like GDPR[142], CCPA [126], and HIPAA [9].
Profiling based on unstructured and noisy data requires significant expertise and is considered too resource-intensive for large-scale privacy breaches [46]. The emergence of LLMs fundamentally alters this landscape. By leveraging their sophisticated understanding and reasoning capabilities, LLMs can automate the inference process, systematically analyzing vast digital footprints to infer sensitive attributes with minimal human intervention. This automation dramatically amplifies the threat, enabling profiling attacks at an unprecedented scale. A growing body of work has demonstrated the feasibility of LLM-driven profiling attacks [49, 158, 165]. In the following, we categorize these attacks along two primary axes: (i) the data modality they target and (ii) their level of automation.
@@
  • Profiling from Textual Activities. Early LLM-based profiling attack [158] assumes that an adversary can access and scrape the public activities (e.g., posts and comments) of a pseudonymous user from the Internet. The adversary then instructs LLMs with prompts to infer predefined sensitive attributes (i.e., eight types of PIIs), within these textual activities. The results showed that powerful models like GPT-4 [127] can achieve performance comparable to human analysts, even when the humans have the advantage of accessing additional contextual information, which LLMs do not have.
  • Profiling from Visual Activities. With the rise of Vision-Language Models (VLMs) [128, 162], research has expanded to include profiling from images and videos, which are ubiquitous on social media platforms like TikTok and Instagram. Specifically, one study [165] designed carefully crafted prompts using chain-of-thought reasoning [174] and automated zooming techniques to direct VLMs to focus on potentially sensitive details in the photos, thus enhancing privacy-infringing inferences. Another significant privacy risk arises from directly inferring a user’s possible geo-location from their pictures [188, 80, 73, 102]. Research has shown that VLMs can outperform even the professional human players in GeoGuessr [2], which raises serious concerns regarding geographic privacy. However, these models are not infallible; they often exhibit significant regional biases, such as a tendency to over-predict well-known landmarks or locations heavily represented in their training data [73].
@@
  • Semi-Automated Profiling. The majority of current research falls into this category, where the core inference task is automated, but significant human effort is still required for data preparation and defining attack objectives. These systems are powerful in controlled settings but face two major limitations in real-world scenarios: (i) Reliance on curated data. Many studies [165, 158, 188, 104] focus on clean, curated textual or image data that is deliberately designed to contain sensitive information, allowing LLMs and VLMs to infer personal attributes. However, in real-world scenarios, user activities are typically noisy and may not be directly related to personal attributes. As a result, the performance of these semi-automated methods would likely degrade significantly when faced with raw, unfiltered activities. (ii) Predefined attribute targets. These attacks are typically configured to search for a fixed set of sensitive attributes (e.g., age, gender, location), which assumes the adversary already knows what to profile from users. However, in the real world, adversaries do not always know what sensitive attributes are present in a user’s activities. This lack of predefined knowledge prevents the attacker from designing specific strategies to target particular attributes, further limiting the applicability of such attacks.
@@
  • Fully-Automated Profiling. To address the limitations of previous approaches, recent work has focused on developing end-to-end automated profiling systems. One example is AutoProfiler [49], an agent-based profiling framework that automatically scrapes, collects, and analyzes potentially sensitive activities from raw, noisy user data. By coordinating with four specialized LLM agents, AutoProfiler fully automates the process of inferring sensitive attributes. This eliminates the need for background knowledge or profiling expertise, making it highly scalable and suitable for deployment on web-scale platforms. Despite its weaker assumptions, the results show that the inferred attributes extend beyond PIIs, uncovering significant amounts of sensitive information. The move toward full automation has profound implications. It means that adversaries no longer need specialized expertise or prior knowledge to launch sophisticated, large-scale profiling attacks.
@@
Privacy Implications. Automated profiling inference can result in serious privacy breaches. One of the most well-known risks is de-anonymization [120, 121]. Study [49] shows that some Reddit users can be de-anonymized by inferring personal attributes from their public activities and comparing these with publicly available profiles, such as LinkedIn. The risk of de-anonymization increases when adversaries gain access to multiple profile databases or cross-reference a user’s activities to construct more comprehensive profiles. In addition, sensitive information extracted from these online activities can also be exploited for severe cybercrimes like doxing and cyberbullying. We refer to [46, 49] for a deeper discussion of the consequences of exposing sensitive personal data.
@@
Growing Threats. Existing attacks exploit the in-context learning (ICL) capability of off-the-shelf LLMs to perform profiling tasks. While this approach is highly efficient and accessible, its performance could be suboptimal, as these models are not specifically designed for profiling. For example, studies show that even state-of-the-art VLMs are outperformed in geo-location identification tasks by PIGEON [64], an image model purpose-built for geolocation. This trend suggests that adversaries may design specialized profiling models that surpass generic LLMs, thereby enhancing attack effectiveness and posing even more severe privacy risks.
@@
Challenges in Evaluation. While various methodologies have been proposed to assess the profiling abilities of LLMs [183, 49, 73, 158], there is still no widely acknowledged benchmark to comprehensively evaluate the associated privacy risks. This issue partially stems from a fundamental ethical dilemma: creating a robust benchmark would require a large dataset of real users’ activities with labeled, sensitive, ground-truth attributes. To address this, some researchers have proposed using synthetic datasets generated by LLM agents [192, 79]. However, the behaviors and data produced by these agents may not accurately reflect the complexities of real human activity, limiting their validity and reliability [49]. In addition, evaluation becomes more complex for fully automated systems that perform open-ended inference without predefined attribute targets. Forcing the model to choose from a candidate list simplifies evaluation but fails to measure the model’s true, unconstrained inference capabilities. Therefore, evaluating the profiling abilities of LLMs remains an open question. Designing effective evaluation approaches is a critical step toward understanding and mitigating these emerging privacy threats.
@@ 5.2 Automated Social Engineering
Automated Investigation. The purpose of this phase is to gather sufficient information about a target to personalize the attack and make it more convincing [123]. Adversaries may directly employ automated profiling strategies (as described in the previous section) to collect personal information. In addition, they may launch proactive information-gathering attempts by manipulating LLM-based chatbots to elicit sensitive details. In such scenarios, a chatbot convinces the user that certain personal information is required to complete a task. Because users often perceive LLMs as helpful assistants, they may willingly provide sensitive details, believing them to be necessary [12, 90, 195, 86]. Attackers can exploit this trust by embedding hidden, privacy-invasive prompts into a chatbot’s behavior [158]. For example, a chatbot tasked with creating a travel itinerary might subtly request additional personal details—such as financial information or contact numbers—under the guise of improving the service. The risk is further amplified in multi-agent LLM systems, where multiple agents collaborate by asking for complementary pieces of information and together constructing a detailed personal profile of the victim [209]. These LLM-based information collection strategies dramatically reduce the cost and time required for reconnaissance while producing highly detailed and actionable intelligence about targets.
@@
Second, LLMs open new avenues for attack by exploiting the growing use of chatbots for emotional and psychological support [207, 72]. In this scenario, attackers deploy malicious chatbots that impersonate trusted friends or companions to establish a deep emotional connection with a victim. The proliferation of third-party platforms like the OpenAI GPT Store [4] and FlowGPT [5] makes it easy to distribute these deceptive chatbots to a wide audience. Once an emotional connection is established, adversaries can manipulate victims into disclosing sensitive information, transferring money under fraudulent pretenses, or even engaging in harmful behaviors [7].

### v2

@@ 1 Introduction
Beyond the risks of training data leakage, privacy threats also emerge from the integration of LLMs into larger, more complex systems, which we refer to as LLM-powered systems. These systems, especially those that use LLMs as decision-making engines in agent-based applications [27, 137], introduce new vulnerabilities and expand the potential attack surface for privacy violations. For instance, a user may share personal information with an LLM-based chatbot in order to receive personalized responses or suggestions. However, this information could be exfiltrated through side channels [14] or unintentionally disclosed by the model itself [iclr24beyond]. Such risks are not inherent to the LLM alone but emerge from the architecture of interactions between users, models, and other system components. As LLM-powered applications become increasingly widespread in both daily life and professional domains, these privacy risks become more prominent and urgent.
A third category of privacy threats arises from the advanced reasoning and autonomous decision-making capabilities of LLMs, which create new opportunities for malicious exploitation. These capabilities enable adversaries to automate sophisticated attacks at unprecedented scale and speed, substantially lowering the barrier to entry for cyberattacks. For instance, an attacker could instruct an LLM to infer sensitive attributes, such as a user’s demographics, from their public online posts, leading to de-anonymization and other severe cybercrimes [arxiv25autoprofiler, 79]. Similarly, LLMs can be leveraged to launch large-scale, highly personalized social engineering campaigns, resulting in significant financial and societal consequences [82].
@@ 3.1 Membership Inference Attack
Closely connected with Differential Privacy (DP) [dwork_dp, ccs13membership], Membership Inference Attack (MIA) [121] has become a widely adopted approach for privacy auditing of machine learning (ML) models [94, 127]. MIA assesses how much a trained ML model reveals about its training data by determining whether specific query instances (or documents) were included in the dataset. In the context of LLMs, membership inference attacks have been applied at different stages to explore the privacy risk.
@@
Potential Mitigation Strategies. Machine learning with differential privacy (e.g., DP-SGD [ccs16dpsgd] and PATE [iclr17pate]) is an effective defense mechanism against privacy attacks, including MIAs. Studies [76, 105, 158] have applied DP-SGD to fine-tune LLMs on sensitive domains, which may degrade model utility under reasonable privacy budgets. For in-context learning scenarios, various DP-based techniques have been proposed, such as differential private prompt tuning [37, 55] and differential private synthetic text generation [147, 130]. In addition to the theoretical privacy guarantees that DP offers, many empirical privacy methods [164, 56] have shown effectiveness against MIAs. For example, LoRA [56], a widely used efficient fine-tuning method for LLMs, demonstrates better privacy preservation than full fine-tuning. However, these empirical defenses may be compromised when facing stronger MIAs [50], and there remains a notable trade-off between privacy and utility, especially for pre-trained LLMs.
@@ 4 Privacy Risks in LLM-Powered Systems
The integration of LLMs as core components in larger, complex systems introduces new vectors for privacy risks. For example, LLM-based chatbots, such as ChatGPT [98] and Claude [5], have become the primary interface through which users interact with LLMs. In these interactions, users often share personal narratives and sensitive details to seek advice or obtain personalized responses. This turns their conversation histories into a rich repository of private information, including personal preferences, habits, and even users’ secrets [iclr24contextual, 9, 66]. Like other computer systems, these applications are vulnerable to side-channel attacks, where adversaries exploit indirect information leaks from the system to steal data. Furthermore, the unique features of LLM-based applications, such as reasoning and memory mechanisms, provide additional attack surfaces. Below, we detail two prominent threats associated with LLM-powered systems: side channel attacks and information exfiltration.
@@ 4.2 Information Exfiltration
  • Unintended Disclosure. LLMs often lack awareness of privacy norms and the contextual boundaries of information flows [97]. As a result, they may inadvertently disclose sensitive information to inappropriate recipients. An LLM agent involved in a multi-round conversation may unintentionally repeat or expose previously shared user information, even when it is contextually irrelevant [90, das2025disclosureauditsllmagents, 117, 169]. For instance, it is undesirable for an LLM assistant to reveal that “John is talking to a few companies about switching jobs” while drafting an email to John’s current manager, particularly without his consent. This risk increases when LLMs are tasked with complex operations that involve integrating multiple sources of user data, such as combining financial, location, and preference information for personalized recommendations [117]. LLMs struggle to track which information is appropriate to share, which makes these disclosures particularly insidious.
@@
  • Leakge via Share Link. Users may share their conversations with a chatbot (e.g., ChatGPT) through a built-in “Share” button on commercial chatbot platforms, allowing only those with the link to access the conversation. However, it has been shown [sharegpt] that these links can be discovered through search engines (e.g., Google Search). This can unintentionally expose users’ conversation histories to the public. Furthermore, deleting a conversation from your ChatGPT history does not remove the public share link or prevent it from appearing in search engine results. Recognizing the potential for privacy breaches, OpenAI has addressed this issue by providing users with the option to control whether their chats are visible in search engine results, offering more control over users’ privacy.
@@
Vulnerability Detection and Mitigation Strategies. To counter the threat of information exfiltration, research efforts are focused on both detecting vulnerabilities and developing active defenses. On the detection front, some works measure an LLM’s capacity for privacy reasoning in ambiguous contexts to identify risks of unintended information disclosure [157, 118]. Another approach [7, zharmagambetov2025agentdam] examines whether agents interacting with web interfaces adhere to the principle of data minimization, introducing benchmarks to systematically evaluate their compliance. In parallel, other efforts aim to build direct defenses against malicious attacks. This includes designing robust countermeasures for prompt injection attacks, which are a primary vector for information exfiltration [31, 136]. Despite these advancements, a comprehensive mitigation strategy capable of defending against the full spectrum of emerging exfiltration threats remains an open challenge [52, 165].
@@ 5.1 Automated Profile Inference
Individuals constantly generate digital footprints through their online activities, encompassing activities from social media comments and posts to shared photos and videos. While some of this data is inherently private (e.g., browse history), a vast amount of these activities (e.g., posts and comments) is publicly accessible. However, the public availability of this information does not eliminate privacy risks. An adversary can aggregate these seemingly innocuous public activities to construct a detailed personal profile, a process known as profiling [profiling, profiling_book]. For instance, analyzing a Reddit user’s most frequented subreddits could reveal their hobbies, while geotags in posted images could disclose their travel patterns or home location. Profiling is widely recognized as a privacy violation by legitimate privacy frameworks like GDPR[gdpr], CCPA [ccpa], and HIPAA [hippa].
Profiling based on unstructured and noisy data requires significant expertise and is considered too resource-intensive for large-scale privacy breaches [16doxing]. The emergence of LLMs fundamentally alters this landscape. By leveraging their sophisticated understanding and reasoning capabilities, LLMs can automate the inference process, systematically analyzing vast digital footprints to infer sensitive attributes with minimal human intervention. This automation dramatically amplifies the threat, enabling profiling attacks at an unprecedented scale. A growing body of work has demonstrated the feasibility of LLM-driven profiling attacks [arxiv25autoprofiler, iclr24beyond, nips24imageinfer]. In the following, we categorize these attacks along two primary axes: (i) the data modality they target and (ii) their level of automation.
@@
  • Profiling from Textual Activities. Early LLM-based profiling attack [iclr24beyond] assumes that an adversary can access and scrape the public activities (e.g., posts and comments) of a pseudonymous user from the Internet. The adversary then instructs LLMs with prompts to infer predefined sensitive attributes (i.e., eight types of PIIs), within these textual activities. The results showed that powerful models like GPT-4 [98] can achieve performance comparable to human analysts, even when the humans have the advantage of accessing additional contextual information, which LLMs do not have.
  • Profiling from Visual Activities. With the rise of Vision-Language Models (VLMs) [99, 131], research has expanded to include profiling from images and videos, which are ubiquitous on social media platforms like TikTok and Instagram. Specifically, one study [nips24imageinfer] designed carefully crafted prompts using chain-of-thought reasoning [142] and automated zooming techniques to direct VLMs to focus on potentially sensitive details in the photos, thus enhancing privacy-infringing inferences. Another significant privacy risk arises from directly inferring a user’s possible geo-location from their pictures [geolocater, arxiv25evaluategeoinfer, arxiv25vlmgeoinfer, arxiv25imageprofiler]. Research has shown that VLMs can outperform even the professional human players in GeoGuessr [geoguessr], which raises serious concerns regarding geographic privacy. However, these models are not infallible; they often exhibit significant regional biases, such as a tendency to over-predict well-known landmarks or locations heavily represented in their training data [arxiv25vlmgeoinfer].
@@
  • Semi-Automated Profiling. The majority of current research falls into this category, where the core inference task is automated, but significant human effort is still required for data preparation and defining attack objectives. These systems are powerful in controlled settings but face two major limitations in real-world scenarios: (i) Reliance on curated data. Many studies [nips24imageinfer, iclr24beyond, geolocater, usenix25piiextract] focus on clean, curated textual or image data that is deliberately designed to contain sensitive information, allowing LLMs and VLMs to infer personal attributes. However, in real-world scenarios, user activities are typically noisy and may not be directly related to personal attributes. As a result, the performance of these semi-automated methods would likely degrade significantly when faced with raw, unfiltered activities. (ii) Predefined attribute targets. These attacks are typically configured to search for a fixed set of sensitive attributes (e.g., age, gender, location), which assumes the adversary already knows what to profile from users. However, in the real world, adversaries do not always know what sensitive attributes are present in a user’s activities. This lack of predefined knowledge prevents the attacker from designing specific strategies to target particular attributes, further limiting the applicability of such attacks.
@@
  • Fully-Automated Profiling. To address the limitations of previous approaches, recent work has focused on developing end-to-end automated profiling systems. One example is AutoProfiler [arxiv25autoprofiler], an agent-based profiling framework that automatically scrapes, collects, and analyzes potentially sensitive activities from raw, noisy user data. By coordinating with four specialized LLM agents, AutoProfiler fully automates the process of inferring sensitive attributes. This eliminates the need for background knowledge or profiling expertise, making it highly scalable and suitable for deployment on web-scale platforms. Despite its weaker assumptions, the results show that the inferred attributes extend beyond PIIs, uncovering significant amounts of sensitive information. The move toward full automation has profound implications. It means that adversaries no longer need specialized expertise or prior knowledge to launch sophisticated, large-scale profiling attacks.
@@
Privacy Implications. Automated profiling inference can result in serious privacy breaches. One of the most well-known risks is de-anonymization [sp08netflix, sp09network]. Study [arxiv25autoprofiler] shows that some Reddit users can be de-anonymized by inferring personal attributes from their public activities and comparing these with publicly available profiles, such as LinkedIn. The risk of de-anonymization increases when adversaries gain access to multiple profile databases or cross-reference a user’s activities to construct more comprehensive profiles. In addition, sensitive information extracted from these online activities can also be exploited for severe cybercrimes like doxing and cyberbullying. We refer to [16doxing, arxiv25autoprofiler] for a deeper discussion of the consequences of exposing sensitive personal data.
@@
Growing Threats. Existing attacks exploit the in-context learning (ICL) capability of off-the-shelf LLMs to perform profiling tasks. While this approach is highly efficient and accessible, its performance could be suboptimal, as these models are not specifically designed for profiling. For example, studies show that even state-of-the-art VLMs are outperformed in geo-location identification tasks by PIGEON [cvpr24pigeon], an image model purpose-built for geolocation. This trend suggests that adversaries may design specialized profiling models that surpass generic LLMs, thereby enhancing attack effectiveness and posing even more severe privacy risks.
@@
Challenges in Evaluation. While various methodologies have been proposed to assess the profiling abilities of LLMs [151, arxiv25autoprofiler, arxiv25vlmgeoinfer, iclr24beyond], there is still no widely acknowledged benchmark to comprehensively evaluate the associated privacy risks. This issue partially stems from a fundamental ethical dilemma: creating a robust benchmark would require a large dataset of real users’ activities with labeled, sensitive, ground-truth attributes. To address this, some researchers have proposed using synthetic datasets generated by LLM agents [nips24pai, arxiv25synpii]. However, the behaviors and data produced by these agents may not accurately reflect the complexities of real human activity, limiting their validity and reliability [arxiv25autoprofiler]. In addition, evaluation becomes more complex for fully automated systems that perform open-ended inference without predefined attribute targets. Forcing the model to choose from a candidate list simplifies evaluation but fails to measure the model’s true, unconstrained inference capabilities. Therefore, evaluating the profiling abilities of LLMs remains an open question. Designing effective evaluation approaches is a critical step toward understanding and mitigating these emerging privacy threats.
@@ 5.2 Automated Social Engineering
Automated Investigation. The purpose of this phase is to gather sufficient information about a target to personalize the attack and make it more convincing [96]. Adversaries may directly employ automated profiling strategies (as described in the previous section) to collect personal information. In addition, they may launch proactive information-gathering attempts by manipulating LLM-based chatbots to elicit sensitive details. In such scenarios, a chatbot convinces the user that certain personal information is required to complete a task. Because users often perceive LLMs as helpful assistants, they may willingly provide sensitive details, believing them to be necessary [sp25chatbot, usenix25llmmental, arxiv25discloursedefense, chi25measuredisclourse]. Attackers can exploit this trust by embedding hidden, privacy-invasive prompts into a chatbot’s behavior [iclr24beyond]. For example, a chatbot tasked with creating a travel itinerary might subtly request additional personal details—such as financial information or contact numbers—under the guise of improving the service. The risk is further amplified in multi-agent LLM systems, where multiple agents collaborate by asking for complementary pieces of information and together constructing a detailed personal profile of the victim [25aidevil]. These LLM-based information collection strategies dramatically reduce the cost and time required for reconnaissance while producing highly detailed and actionable intelligence about targets.
@@
Second, LLMs open new avenues for attack by exploiting the growing use of chatbots for emotional and psychological support [chi25llmemotional, nature25mental]. In this scenario, attackers deploy malicious chatbots that impersonate trusted friends or companions to establish a deep emotional connection with a victim. The proliferation of third-party platforms like the OpenAI GPT Store [gptstore] and FlowGPT [flowgpt] makes it easy to distribute these deceptive chatbots to a wide audience. Once an emotional connection is established, adversaries can manipulate victims into disclosing sensitive information, transferring money under fraudulent pretenses, or even engaging in harmful behaviors [suicidegpt].